OpenStudy (anonymous):

I am trying to help my daughter with summer algebra and can't remember what steps to solve (a to the 5th times bsquared all the the 3rd divided by a to the 7th and b to the 3rd. We know the answer is a to the 8th times b 3rd. Would love help!

OpenStudy (anonymous):

Remember when you raise a power to a power, you multiply exponents: \[\large (a^5b^2)^3 = a^{5\cdot 3}b^{2\cdot 3} = a^{15}b^{6}\] And when you divide powers of the same base you subtract their exponents. \[\large \frac{ a^{15}b^{6}}{a^7b^3} = a^{15-7}b^{6-3} = a^8b^3\]

OpenStudy (amistre64):

a^5 is notation for "a to the 5th" b^2 is notation for "bsquared" (a^5 b^2)^3 is likewise easier to read mathematically .. just saying :)

OpenStudy (amistre64):

at times people omit the "^" part and its just as sensible for the most part